Constants Part A A parallel-plate capacitor with only air between the plates is charged by connecting it to a battery The capacitor is then disconnected from the battery without any of the charge leaving the plates A voltmeter reads 49 0 V when placed across the capacitor When a dielectric is inserted between the plates, completely filling the space, the voltmeter reads 14.5 V What is the dielectric constant of this material? Part B What will the voltmeter read il the dielectric is out so it tils only one third of the space between the plates? Pearsorn
Answer #1

Given,

V = 49 V ; V' = 14.5 V

We know that

Q = CV => C = Q/V

C' = Q/V'

C' = k C

Also, Q' = Q

k Q/V = Q/V'

k = V/V' = 49/14.5 = 3.38

Hence, k = 3.38

b)C1 = e0 2A/3d

C2 = ke0 A/3d

Ceq = C(2/3 + 1/3k)

V = Q/C

V = Q/C(2/3 + 1/3k)

V = 49/(2/3 + 1/3 x 3.38) = 27.32 V

Hence, V = 27.32 V